OpenStudy (anonymous):

convert to fractional notation and simplify your answer 2 1/7

OpenStudy (anonymous):

15/7

OpenStudy (anonymous):

u always multiply the # with the denominator and add it to the munerator!!

OpenStudy (anonymous):

so 2 x 7 = 14 + 1 =15/7